3) Yellow pine has a mass density of 0.55 g/cm3. Suppose you have a wooden yellow pine cylinder with a volume of 8.3 cm3 .

A) What is the mass of the cylinder?

B) What is the weight of the cylinder (the force due to gravity)?

C) If you immerse the cylinder in water, what will the buoyant force on it be?

D) What do your results from parts B and C imply?
Answer #1

biven, mais density (3) = 0.55g/cm3 Volume (V) = 8.3cm3 4 mars (m) = A. use density = volume m = 8v = 0.559/cm² x 8 [m= 4.6 g